OpenStudy (anonymous):

If a coat that originally cost $115 was marked down 20%, how much would be marked off the price? A. $22 B. $23 C. $20 D. $24

OpenStudy (mertsj):

What percent is taken off?

OpenStudy (mertsj):

gage???? Are you there?

Directrix (directrix):

B. $23 $23 would be marked off the original price .20(115) =23

OpenStudy (anonymous):

\[\LARGE 20\% \quad \longrightarrow \quad \frac{20}{100}\] \[\LARGE 115(1-20\%)=115-115\cdot \frac{20}{100}=92\] so... \[\LARGE 115-92=23 \]

OpenStudy (turingtest):

@Directrix please try to provide a little more explanation. This is a teaching site after all.

OpenStudy (anonymous):

@TuringTest I agree !

Directrix (directrix):

@TuringTest --> My work is shown. File a CoC. Hollering at me on this site helps no one.